OpenStudy (goformit100):

A, B, C, D and E play a game of cards. A says to B, "If you give me three cards, you will have as many as E has and if I give you three cards, you will have as many as D has." A and B together have 10 cards more than what D and E together have. If B has two cards more than what C has and the total number of cards be 133, how many cards does B have ? A. 22 B. 23 C. 25 D. 35

OpenStudy (goformit100):

@LoveYou*69

OpenStudy (goformit100):

@charlotte123

OpenStudy (goformit100):

@allie_bear22 @lillybeth123

OpenStudy (anonymous):

sorry, idk, :(

OpenStudy (anonymous):

25

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Clearly, we have : B-3 = E B + 3 = D A+B = D + E+10 B = C + 2 A+B + C + D + E= 133 solving the above equations you'll get the number of cards B have